# Depth Pro: Sharp Monocular Metric Depth in Less Than a Second

### v1

@@ 1 Introduction
Depth Pro is enabled by a number of technical contributions. First, we design an efficient multi-scale ViT-based architecture for capturing the global image context while also adhering to fine structures at high resolution. Second, we derive a new set of metrics that enable leveraging highly accurate matting datasets for quantifying the accuracy of boundary tracing in evaluating monocular depth maps. Third, we devise a set of loss functions and a training curriculum that promote sharp depth estimates while training on real-world datasets that provide coarse and inaccurate supervision around boundaries, along with synthetic datasets that offer accurate pixelwise ground truth but limited realism. Fourth, we contribute zero-shot focal length estimation from a single image that dramatically outperforms the prior state of the art.
@@ 2 Related work
Zero-shot depth estimation. MegaDepth (Li & Snavely, 2018) demonstrated that training on a diverse dataset allows generalizing monocular depth prediction beyond a specific domain. MiDaS (Ranftl et al., 2022) advanced this idea by training on a large mix of diverse datasets with a scale-and-shift-invariant loss. Follow-up works applied this recipe to transformer-based architectures (Ranftl et al., 2021; Birkl et al., 2023) and further expanded the set of feasible datasets through self-supervision (Spencer et al., 2023; Yang et al., 2024a). A line of work uses self-supervision to learn from unlabeled image and video data (Petrovai & Nedevschi, 2022; Yang et al., 2024a). A number of recent approaches (Ke et al., 2024; Gui et al., 2024) harness diffusion models to synthesize relative depth maps. Although some of these methods demonstrated excellent generalization, their predictions are ambiguous in scale and shift, which precludes downstream applications that require accurate shapes, sizes, or distances.
@@
Sharp occluding contours. SharpNet (Ramamonjisoa & Lepetit, 2019) incorporates normals and occluding contour constraints, but requires additional contour and normal supervision during training. BoostingDepth (Miangoleh et al., 2021) obtains detailed predictions from a low-resolution network by applying it independently to image patches. Since the patches lack global context, BoostingDepth fuses them through a sophisticated multi-step pipeline. PatchFusion (Li et al., 2024a) refines this concept through image-adaptive patch sampling and tailored modules that enable end-to-end training. A recent line of work leverages diffusion priors to enhance the sharpness of occlusion boundaries (Gui et al., 2024; Ke et al., 2024). These approaches predominantly focus on predicting relative (rather than metric) depth. We propose a simpler architecture without task-specific modules or diffusion priors and demonstrate that even sharper and more accurate results can be obtained while producing metric depth maps and reducing runtime by more than two orders of magnitude.
@@
Multi-scale vision transformers. Vision transformers (ViTs) have emerged as the dominant general-purpose architecture for perception tasks, but operate at low resolution (Dosovitskiy et al., 2021). Naïvely scaling the architecture to higher resolutions is prohibitive due to the computational complexity. Several works identified the attention layers as the main obstacle to scaling up ViT and have proposed alternatives (Zhu et al., 2021; Liu et al., 2021; Li et al., 2022c; Chu et al., 2021; Liu et al., 2022a; 2023; Cai et al., 2023; Jaegle et al., 2022).
Another line of work modified the ViT architecture to produce a hierarchy of features (Fan et al., 2021; Xie et al., 2021; Yuan et al., 2021; Ranftl et al., 2021; Chen et al., 2021; Lee et al., 2022).
Rather than modifying the ViT architecture, which requires computationally expensive retraining, we propose a network architecture that applies a plain ViT backbone at multiple scales and fuses predictions into a single high-resolution output. This architecture benefits from ongoing improvements in ViT pretraining, as new variants can be easily swapped in (Oquab et al., 2024; Peng et al., 2022b; Sun et al., 2023).
@@ 3.1 Network
The key idea of our architecture is to apply plain ViT encoders (Dosovitskiy et al., 2021) on patches extracted at multiple scales and fuse the patch predictions into a single high-resolution dense prediction in an end-to-end trainable model. Fig. 3 illustrates the architecture. For predicting depth, we employ two ViT encoders, a patch encoder and an image encoder. The patch encoder is applied on patches extracted at multiple scales. Intuitively, this may allow learning representations that are scale-invariant as weights are shared across scales. The image encoder anchors the patch predictions in a global context. It is applied to the whole input image, downsampled to the base input resolution of the chosen encoder backbone (in our case 384384).
The whole network operates at a fixed resolution of 15361536, which was chosen as a multiple of the ViT’s 384384. This guarantees a sufficiently large receptive field and constant runtimes for any image while preventing out-of-memory errors (which we repeatedly observed for variable-resolution approaches on large images). Confirming this design choice, the results we report in Sec. 4 and Tab. 5 demonstrate that Depth Pro is consistently orders of magnitude faster than variable-resolution approaches while being more accurate and producing sharper boundaries. A key benefit of assembling our architecture from plain ViT encoders over custom encoders is the abundance of pretrained ViT-based backbones that can be harnessed (Oquab et al., 2024; Peng et al., 2022b; Sun et al., 2023). We evaluate several pretrained backbones and compare our architecture to other high-resolution architectures in the appendices (Tab. 8 and Sec. B.2).
After downsampling to 15361536, the input image is split into patches of . For the two finest scales, we let patches overlap to avoid seams. At each scale, the patches are fed into the patch encoder, which produces a feature tensor at resolution per input patch (features 3 – 6 in Fig. 3). At the finest scale we further extract intermediate features (features 1 & 2 in Fig. 3) to capture finer-grained details. We merge the feature patches into maps, which are fed into the decoder module, which resembles the DPT decoder (Ranftl et al., 2021).
@@ 3.2 Sharp monocular depth estimation
We train with several objectives, all based on canonical inverse depth, because this prioritizes areas close to the camera over farther areas or the whole scene, and thus supports visual quality in applications such as novel view synthesis. Let be the ground-truth canonical inverse depth. For all metric datasets we compute the mean absolute error (, Eq. 1) per pixel , and discard pixels with an error in the top per image for real-world (as opposed to synthetic) datasets:
@@
Based on these observations, we design a two-stage training curriculum. In the first stage, we aim to learn robust features that allow the network to generalize across domains. To that end, we train on a mix of all labeled training sets. Specifically, we minimize on metric datasets and its normalized version on non-metric datasets. is chosen for its robustness in handling potentially corrupted real-world ground truth. To steer the network towards sharp boundaries, we aim to also supervise on gradients of the predictions. Done naïvely, however, this can hinder optimization and slow down convergence. We found that a scale-and-shift-invariant loss on gradients, applied only to synthetic datasets, worked best. Controlled experiments are reported in the appendices.
@@
To penalize blurry edges, we suppress non-maximum values of within the valid bounds of connected components. For additional experiments and qualitative results we refer to the appendices.
@@ 3.3 Focal length estimation
To handle images that may have inaccurate or missing EXIF metadata, we supplement our network with a focal length estimation head. A small convolutional head ingests frozen features from the depth estimation network and task-specific features from a separate ViT image encoder to predict the horizontal angular field-of-view. We use as the training loss. We train the focal length head and the ViT encoder after the depth estimation training. Separating the focal length training has several benefits over joint training with the depth network. It avoids the necessity of balancing the depth and focal length training objectives. It also allows training the focal length head on a different set of datasets, excluding some narrow-domain single-camera datasets that are used in training the depth estimation network, and adding large-scale image datasets that provide focal length supervision but no depth supervision. Further details are provided in the appendices.
@@ 4 Experiments
Zero-shot boundaries. Tab. 2 summarizes the evaluation of boundary accuracy for Depth Pro and a number of baselines. This evaluation is conducted in a zero-shot setting: models are only evaluated on datasets that were not seen during training. Since our boundary metrics are scale-invariant, our baselines here also include methods that only predict relative (rather than absolute metric) depth. Our absolute baselines include Metric3D (Yin et al., 2023), Metric3D v2 (‘giant’ model) (Hu et al., 2024), PatchFusion (Li et al., 2024a), UniDepth (Piccinelli et al., 2024), and ZoeDepth (Bhat et al., 2023). We also report results for the relative variants of Depth Anything v1 & v2 (Yang et al., 2024a; b) because they yield sharper boundaries than their metric counterparts. Lastly, we include Marigold (Ke et al., 2024), a recent diffusion-based relative depth model that became popular due to its high-fidelity predictions. We use the boundary metrics introduced in Sec. 3.2, and report the average boundary F1 score for datasets with ground-truth depth, and boundary recall () for datasets with matting or segmentation annotations. For image matting datasets, a pixel is marked as occluding when the value of the alpha matte is above .
@@
We find that Depth Pro produces more accurate boundaries than all baselines on all datasets, by a significant margin. As can be observed in Fig. 1, in the images in Tab. 2, and the additional results in Sec. A, the competitive metric accuracy of the recent Metric3D v2 and Depth Anything v2 models does not imply sharp boundaries. Depth Pro has a consistently higher recall for thin structures like hair and fur and yields sharper boundaries. This is also true in comparison to the diffusion-based Marigold, which leverages a prior trained on billions of real-word images, as well as PatchFusion, which operates at variable resolution. Note that the runtime of Depth Pro is orders of magnitude faster than Marigold and PatchFusion (see Fig. 2 & Tab. 5). Fig. 4 demonstrates the benefits of sharp boundary prediction for novel view synthesis from a single image.
@@
Focal length estimation. Previous work (Piccinelli et al., 2024; Kocabas et al., 2021; Baradad & Torralba, 2020) does not provide comprehensive systematic evaluations of focal length estimators on in-the-wild images. To address this, we curated a zero-shot test dataset. To this end, we selected diverse datasets with intact EXIF data, enabling reliable assessment of focal length estimation accuracy. FiveK (Bychkovsky et al., 2011), DDDP (Abuolaim & Brown, 2020), and RAISE (Dang-Nguyen et al., 2015) contribute professional-grade photographs taken with SLR cameras. SPAQ (Fang et al., 2020) provides casual photographs from mobile phones. PPR10K (Liang et al., 2021) provides high-quality portrait images. Finally, ZOOM (Zhang et al., 2019) includes sets of scenes captured at various optical zoom levels.
@@
Tab. 3 compares Depth Pro against state-of-the-art focal length estimators and shows the percentage of images with relative estimation error under 25% and 50%, respectively. Depth Pro is the most accurate across all datasets. For example, on PPR10K, a dataset of human portraits, our method leads with 64.6% of the images having a focal length error below 25%, while the second-best method, SPEC, only achieves 34.6% on this metric. We attribute this superior performance to our network design and training protocol, which decouple training of the focal length estimator from the depth network, enabling us to use different training sets for these two tasks. Further controlled experiments are reported in the appendices.
@@ 5 Conclusion & limitations
Depth Pro produces high-resolution metric depth maps with high-frequency detail at sub-second runtimes. Our model achieves state-of-the-art zero-shot metric depth estimation accuracy without requiring metadata such as camera intrinsics, and traces out occlusion boundaries in unprecedented detail, facilitating applications such as novel view synthesis from single images ‘in the wild’. While Depth Pro outperforms prior work along multiple dimensions, it is not without limitations. For example, the model is limited in dealing with translucent surfaces and volumetric scattering, where the definition of single pixel depth is ill-posed and ambiguous.
@@ A.1 Qualitative results
We provide additional qualitative results of Depth Pro, Marigold (Ke et al., 2024), Metric3D v2 (Hu et al., 2024), and Depth Anything v2 (Yang et al., 2024b) on in-the-wild images from AM-2k (Li et al., 2022a), DIS-5k (Qin et al., 2022), and Unsplash111https://www.unsplash.com in Fig. 5, Fig. 6, and Fig. 7. Fine details are repeatedly missed by Metric3D v2 and Depth Anything v2. Marigold reproduces finer details than Metric3D v2 and Depth Anything v2, but commonly yields noisy predictions.
@@ Appendix B Controlled Experiments
We conduct several controlled experiments to investigate the impact of various components and design decisions in Depth Pro. Specifically, we aim to assess the contribution of key components in the network architecture (Sec. B.1), training objective (Sec. B.3), training curriculum (Sec. B.4), and the focal length estimation head (Sec. B.5).
B.1 Network backbone
We begin by evaluating various candidate image encoder backbones within our network architecture. To assess their performance, we conduct a comparative analysis utilizing off-the-shelf models available from the TIMM library (Wightman, 2019). Using the pretrained weights, we train each backbone at 384 384 resolution across five RGB-D datasets (Keystone, HRWSI, RedWeb, TartanAir, and Hypersim) and evaluate their performance in terms of metric depth accuracy across multiple datasets, including Booster, Hypersim, Middlebury, and NYUv2, utilizing metrics such as for affine-invariant depth and for metric depth in Tab. 7. We find that ViT-L DINOv2 (Oquab et al., 2024) outperforms all other backbones by a significant margin and conclude that the combination of backbone and pretraining strategy considerably affects downstream performance. Following this analysis, we pick ViT-L DINOv2 for both our encoder backbones.
B.2 High-resolution alternatives
We further evaluate alternative high-resolution 15361536 network structures and different pre-trained weights (Tab. 8). To do this, we test generalization accuracy by training on a train split of some datasets and testing on a val or test split of other datasets, following the Stage 1 protocol for all models in accordance with Tab. 14 and Tab. 15. All ViT models use a patch size of 1616. For weights pretrained with a patch size of 1414 we apply bicubic interpolation to the weights of the convolutional patch embedding layer and scale these weights inversely to the number of pixels (i.e., the weights are reduced by a factor of 1.3). All ViT models use resolution 15361536, for this we apply bicubic interpolation to positional embeddings prior to training. The Depth Pro approach in all cases uses ViT with resolution 384384 and patch size 1616 for both the patch encoder and the image encoder. SWINv2 and convolutional models are pretrained on ImageNet (Deng et al., 2009). Other models use different pretraining approaches described in their papers: CLIP (Radford et al., 2021), MAE (He et al., 2022), BeiTv2 (Peng et al., 2022b), and DINOv2 (Oquab et al., 2024). For the Segment Anything model we use publicly available pretrained weights, which were initialized using MAE pretraining (He et al., 2022) and subsequently trained for segmentation as described in their paper (Kirillov et al., 2023).

### v2

@@ 1 Introduction
Depth Pro is enabled by a number of technical contributions. First, we derive a new set of metrics that enable leveraging highly accurate matting datasets for quantifying the accuracy of boundary tracing in evaluating monocular depth maps. We analyze the effect of typical output resolutions and find that a high resolution is necessary but not sufficient to improve boundary accuracy. Second, we design an efficient multi-scale ViT-based architecture for capturing the global image context while also adhering to fine structures at high resolution. Third, we devise a set of loss functions and a training curriculum that promote sharp depth estimates while training on real-world datasets that provide coarse and inaccurate supervision around boundaries, along with synthetic datasets that offer accurate pixelwise ground truth but limited realism. Fourth, we contribute zero-shot focal length estimation from a single image that dramatically outperforms the prior state of the art.
@@ 2 Related work
Zero-shot depth estimation. MegaDepth (Li & Snavely, 2018) demonstrated that training on a diverse dataset allows generalizing monocular depth prediction beyond a specific domain. MiDaS (Ranftl et al., 2022) advanced this idea by training on a large mix of diverse datasets with a scale-and-shift-invariant loss. Follow-up works applied this recipe to transformer-based architectures (Ranftl et al., 2021; Birkl et al., 2023) and further expanded the set of feasible datasets through self-supervision (Spencer et al., 2023; Yang et al., 2024a). A line of work uses self-supervision to learn from unlabeled image and video data (Petrovai & Nedevschi, 2022; Yang et al., 2024a). A number of recent approaches (Ke et al., 2024; Gui et al., 2025) harness diffusion models to synthesize relative depth maps. Although some of these methods demonstrated excellent generalization, their predictions are ambiguous in scale and shift, which precludes downstream applications that require accurate shapes, sizes, or distances.
@@
Sharp occluding contours. SharpNet (Ramamonjisoa & Lepetit, 2019) incorporates normals and occluding contour constraints, but requires additional contour and normal supervision during training. BoostingDepth (Miangoleh et al., 2021) obtains detailed predictions from a low-resolution network by applying it independently to image patches. Since the patches lack global context, BoostingDepth fuses them through a sophisticated multi-step pipeline. PatchFusion (Li et al., 2024a) refines this concept through image-adaptive patch sampling and tailored modules that enable end-to-end training. A recent line of work (Gui et al., 2025; Ke et al., 2024) leverages diffusion priors to enhance the sharpness of occlusion boundaries. These approaches predominantly focus on predicting relative (rather than metric) depth. We propose a simpler architecture without task-specific modules or diffusion priors and demonstrate that even sharper and more accurate results can be obtained while producing metric depth maps and reducing runtime by more than two orders of magnitude.
@@
Multi-scale vision transformers. Vision transformers (ViTs) have emerged as the dominant general-purpose architecture for perception tasks but operate at low resolution (Dosovitskiy et al., 2021). The computational complexity of their attention layers prohibits naïvely scaling them to higher resolutions and several works proposed alternatives (Zhu et al., 2021; Liu et al., 2021; Li et al., 2022c; Chu et al., 2021; Liu et al., 2022a; 2023; Cai et al., 2023; Jaegle et al., 2022). Another line of work modified the ViT architecture to produce a hierarchy of features (Fan et al., 2021; Xie et al., 2021; Yuan et al., 2021; Ranftl et al., 2021; Chen et al., 2021; Lee et al., 2022).
Rather than modifying the ViT architecture, which requires computationally expensive retraining, we propose an architecture that applies a plain ViT backbone at multiple scales and fuses predictions into a single high-resolution output. This design benefits from ongoing improvements in ViT pretraining, as new variants can be easily swapped in (Oquab et al., 2024; Peng et al., 2022b; Sun et al., 2023).
@@ 3.1 Network
The key idea behind our network is to apply plain ViT encoders (Dosovitskiy et al., 2021) on patches extracted at multiple scales and fuse their predictions into a single high-resolution dense depth prediction in an end-to-end trainable model (see Fig. 3). As the patch encoder shares weights across all scales, it may intuitively learn a scale-invariant representation. The image encoder anchors the patch predictions in a global context. It is applied to the whole input image, downsampled to the base input resolution of the chosen encoder backbone (in our case 384384).
The whole network operates at a fixed resolution of 15361536, which was chosen as a multiple of the ViT’s 384384. This guarantees a sufficiently large receptive field and constant runtimes for any image while preventing out-of-memory errors (which we repeatedly observed for variable-resolution approaches on large images). Confirming this design choice, our results in Sec. 4 and Tab. 5 demonstrate that Depth Pro is consistently orders of magnitude faster than variable-resolution approaches while being more accurate and producing sharper boundaries. A key benefit of assembling our architecture from plain ViT encoders over custom encoders is the abundance of pretrained ViT-based backbones that can be harnessed (Oquab et al., 2024; Peng et al., 2022b; Sun et al., 2023). We evaluate several pretrained backbones and compare our architecture to other high-resolution architectures in the appendices (Tab. 9 and Sec. B.3).
After initial downsampling to 15361536, the input image is split into patches of at each scale. For the two finest scales, we let patches overlap to avoid seams, which yields 25 and 9 patches, respectively. In total, we extract 35 patches, concatenate them along the batch dimension to allow efficient batch processing, and feed them to the patch encoder. This yields a feature tensor at resolution per input patch (Features 3 – 6 in Fig. 3). At the finest scale we further extract intermediate features (Features 1 & 2 in Fig. 3) to capture finer-grained details, yielding additional feature patches. We merge the feature patches into maps (detailed in Sec. C.1), which are fed into the decoder module, which resembles the DPT decoder (Ranftl et al., 2021).
@@ 3.2 Sharp monocular depth estimation
We train with several objectives, all based on canonical inverse depth, because this prioritizes areas close to the camera over farther areas or the whole scene, and thus supports visual quality in applications such as novel view synthesis (see Sec. B.4). Let be the ground-truth canonical inverse depth. For all metric datasets we compute the mean absolute error (, Eq. 1) per pixel , and discard pixels with an error in the top per image for real-world (as opposed to synthetic) datasets:
@@
Based on these observations, we design a two-stage training curriculum. In the first stage, we aim to learn robust features that allow the network to generalize across domains. To that end, we train on a mix of all labeled training sets. Specifically, we minimize on metric datasets and its normalized version on non-metric datasets. is chosen for its robustness in handling potentially corrupted real-world ground truth. To steer the network towards sharp boundaries, we aim to also supervise on gradients of the predictions. Done naïvely, however, this can hinder optimization and slow down convergence. We found that a scale-and-shift-invariant loss on gradients, applied only to synthetic datasets, worked best. Controlled experiments are reported in Sec. B.5 of the appendices.
@@
To penalize blurry edges, we suppress non-maximum values of within the valid bounds of connected components. We report additional experiments and qualitative results in Sec. A.4.
@@ 3.3 Focal length estimation
To handle images that may have inaccurate or missing EXIF metadata, we supplement our network with a focal length estimation head. A small convolutional head ingests frozen features from the depth estimation network and task-specific features from a separate ViT image encoder to predict the horizontal angular field-of-view. We use as the training loss. We train the focal length head and the ViT encoder after the depth estimation training. Separating the focal length training has several benefits over joint training with the depth network. It avoids the necessity of balancing the depth and focal length training objectives. It also allows training the focal length head on a different set of datasets, excluding some narrow-domain single-camera datasets that are used in training the depth estimation network, and adding large-scale image datasets that provide focal length supervision but no depth supervision. Further details are provided in Sec. C.2.
@@ 4 Experiments
Zero-shot boundaries. Tab. 2 summarizes the evaluation of boundary accuracy for Depth Pro and several baselines. This evaluation is conducted in a zero-shot setting: models are only evaluated on datasets that were not seen during training. Since our boundary metrics are scale-invariant, our baselines here also include methods that only predict relative (rather than absolute metric) depth. Our absolute baselines include Metric3D (Yin et al., 2023), Metric3D v2 (‘giant’ model) (Hu et al., 2024), PatchFusion (Li et al., 2024a), UniDepth (Piccinelli et al., 2024), and ZoeDepth (Bhat et al., 2023). We also report results for the relative variants of Depth Anything v1 & v2 (Yang et al., 2024a; b) because they yield sharper boundaries than their metric counterparts. Lastly, we include Marigold (Ke et al., 2024), a recent diffusion-based relative depth model that became popular due to its high-fidelity predictions. We use the boundary metrics introduced in Sec. 3.2, and report the average boundary F1 score for datasets with ground-truth depth, and boundary recall () for datasets with matting or segmentation annotations. For image matting datasets, a pixel is marked as occluding when the value of the alpha matte is above .
@@
We find that Depth Pro produces more accurate boundaries than all baselines on all datasets, by a significant margin. As can be observed in Fig. 1, in the images in Tab. 2, and the additional results in Sec. A, the competitive metric accuracy of Metric3D v2 and Depth Anything v2 does not imply sharp boundaries. Depth Pro has a consistently higher recall for thin structures like hair and fur and yields sharper boundaries. This is also true in comparison to the diffusion-based Marigold, as well as PatchFusion, which operates at variable resolution. Note that Depth Pro is orders of magnitude faster than Marigold and PatchFusion (see Fig. 2 & Tab. 5). Fig. 4 demonstrates the benefits of sharp boundary prediction for novel view synthesis from a single image.
@@
Focal length estimation. Prior work (Piccinelli et al., 2024; Kocabas et al., 2021; Baradad & Torralba, 2020) does not provide comprehensive systematic evaluations of focal length estimators on in-the-wild images. To address this, we curated a zero-shot test dataset. To this end, we selected diverse datasets with intact EXIF data, enabling reliable assessment of focal length estimation accuracy. FiveK (Bychkovsky et al., 2011), DDDP (Abuolaim & Brown, 2020), and RAISE (Dang-Nguyen et al., 2015) contribute professional-grade photographs taken with SLR cameras. SPAQ (Fang et al., 2020) provides casual photographs from mobile phones. PPR10K (Liang et al., 2021) provides high-quality portrait images. Finally, ZOOM (Zhang et al., 2019) includes sets of scenes captured at various optical zoom levels.
@@
Tab. 3 compares Depth Pro against state-of-the-art focal length estimators and shows the percentage of images with relative estimation error under 25% and 50%, respectively. Depth Pro is the most accurate across all datasets. For example, on PPR10K, a dataset of human portraits, our method leads with 64.6% of the images having a focal length error below 25%, while the second-best method, SPEC, only achieves 34.6% on this metric. We attribute this superior performance to our network design and training protocol, which decouple training of the focal length estimator from the depth network, enabling us to use different training sets for these two tasks. Further controlled experiments are reported in Sec. B.7 of the appendices.
@@ 5 Conclusion & limitations
Depth Pro produces high-resolution metric depth maps with high-frequency detail at sub-second runtimes. Our model achieves state-of-the-art zero-shot metric depth estimation accuracy without requiring metadata such as camera intrinsics, and traces out occlusion boundaries in unprecedented detail, facilitating applications such as novel view synthesis from single images ‘in the wild’. While Depth Pro outperforms prior work along multiple dimensions, it is not without limitations. For example, the model is limited in dealing with translucent surfaces and volumetric scattering, where the definition of pixelwise depth is ill-posed and ambiguous.
@@ A.1 Qualitative results
We provide additional qualitative results of Depth Pro, Marigold (Ke et al., 2024), Metric3D v2 (Hu et al., 2024), and Depth Anything v2 (Yang et al., 2024b) on in-the-wild images from AM-2k (Li et al., 2022a), DIS-5k (Qin et al., 2022), and Unsplash111https://www.unsplash.com in Fig. 5, Fig. 6, and Fig. 7. Fine details are repeatedly missed by Metric3D v2 and Depth Anything v2. Marigold reproduces finer details than Metric3D v2 and Depth Anything v2 but commonly yields noisy predictions.
@@ Appendix B Controlled Experiments
We conduct several controlled experiments to investigate the impact of various components and design decisions in Depth Pro. Specifically, we aim to assess the contribution of the native output resolution, key components in the network architecture (Sec. B.2), the depth representation, training objectives (Sec. B.5), training curriculum (Sec. B.6), and the focal length estimation head (Sec. B.7).
B.1 Native output resolution
To assess the importance of the native output resolution, we conduct a simple experiment on UnrealStereo4K. We downsample the ground truth depth maps to several common resolutions found in the literature as well the output resolution of Depth Pro. We then upsample again to the original resolution and evaluate depth and boundary metrics. Results are listed in Tab. 7. We find that the native output resolution has a strong effect on boundary accuracy. More specifically, doubling the resolution may improve boundary accuracy by a factor of 3. It is important to note however, that these results represent an upper bound. The zero-shot comparison to the state of the art (Tab. 1 and 2) includes approaches that predict depth at the full input resolution, namely PatchFusion and ZeroDepth. Although PatchFusion and ZeroDepth predict at a higher resolution than e.g., Metric3D v2 or DepthAnything v2, they trail the lower resolution approaches in performance. We conclude that predicting at a high native output resolution is necessary but not sufficient for predicting accurate metric depth with sharp boundaries.
Output resolution | Example | Approximate optimum | ||
Log10 | AbsRel | F1 | ||
15361536 | Depth Pro | 0.019 | 0.004 | 0.311 |
768768 | Marigold (Ke et al., 2024) | 0.048 | 0.010 | 0.131 |
518518 | Depth Anything v2 (Yang et al., 2024a) | 0.084 | 0.016 | 0.065 |
384384 | DPT (Ranftl et al., 2021) | 0.123 | 0.024 | 0.044 |
B.2 Network backbone
To assess the effect of the network architecture, we begin by evaluating various candidate image encoder backbones within our network architecture. To assess their performance, we conduct a comparative analysis utilizing off-the-shelf models available from the TIMM library (Wightman, 2019). Using the pretrained weights, we train each backbone at 384 384 resolution across five RGB-D datasets (Keystone, HRWSI, RedWeb, TartanAir, and Hypersim) and evaluate their performance in terms of metric depth accuracy across multiple datasets, including Booster, Hypersim, Middlebury, and NYUv2, utilizing metrics such as for affine-invariant depth and for metric depth in Tab. 8. We find that ViT-L DINOv2 (Oquab et al., 2024) outperforms all other backbones by a significant margin and conclude that the combination of backbone and pretraining strategy considerably affects downstream performance. Following this analysis, we pick ViT-L DINOv2 for our encoder backbones.